### v1

@@ 3.2 Preventing breakdown by incorporating look-ahead
Hence, if $l_{jj}$ is small after $k<j$ steps then, on subsequent steps, it can only decrease further, not increase. Thus to detect potential B1 breakdown as early as possible, look-ahead can be used whereby, at each major step $k$ of the factorization, the current values of the all remaining diagonal entries $l_{jj}$, $k+1\leq j\leq n$ are updated (using safe operations). If any updated diagonal entry $l_{jj}$ is less than a prescribed tolerance $\tau_{u}$ then B1 breakdown has been detected for column $j$ and the factorization is terminated. It is straightforward to incorporate look-ahead within many IC algorithms (to facilitate this, for some IC variants, it may be necessary to hold a copy of the diagonal entries). Its use is standard if numerical pivoting is incorporated (that is, at step $k$, rows and columns of the remaining matrix are symmetrically interchanged to ensure $l_{kk}$ is larger than the current values of $l_{jj}$ for $j\geq k$) or if a symmetric indefinite variant that employs $2\times 2$ pivots is used [14]. Look-ahead is employed in some well-known fp64 implementations of IC factorizations (e.g., [15, 16]). Algorithm 2 is a modified version of Algorithm 1 that includes both safe checks for breakdown and look-ahead. The safe test for B3 breakdown does not have to be applied to individual scalar entries: but it can be cheaply applied to each column $j\geq k+1$. Once breakdown is detected, the factorization is terminated.
Algorithm 2 Right-looking IC factorization with safe checks for breakdown and look-ahead
@@
Output: Either $flag=0$ and $A\approx L^{T}L$ or $flag>0$ (breakdown detected)
1:Initialize $l_{ij}=a_{ij}$ for all $(i,j)\in{\mathcal{S}}\{L\}$
2:Set $flag=0$
@@
4:for $k=1:n$ donormal-$\triangleright$ Start of $k$-th major step
@@
6:     if  $l_{kk}\geq 1$ or $l_{kk}\geq l_{max}/x_{max}$ thennormal-$\triangleright$ $l_{max}$ is largest off-diagonal entry (2.1)
7:         for $i=k+1:n$ such that $(i,k)\in{\mathcal{S}}\{L\}$ do
@@
9:         end fornormal-$\triangleright$ Column $k$ of $L$ has been computed
@@
11:         $flag=2$ and returnnormal-$\triangleright$ B2 breakdown
@@
13:     for $j=k+1:n$ such that $(j,k)\in{\mathcal{S}}\{L\}$ donormal-$\triangleright$ Update columns $j>k$ of $L$
14:         for $i=j:n$ such that $(i,j)\in{\mathcal{S}}\{L\}$  do
15:              Test column $j$ can be updated safely $\triangleright$ Use Algorithm 2.2 of [6]
@@
18:              if  $i=j$ and $l_{ii}<\tau_{u}$ then
19:                  $flag=1$ and return $\triangleright$ Look-ahead detected B1 breakdown
20:              end if
@@
22:     end for
23:end for
To illustrate the usefulness of look-ahead, consider the following matrix
@@
It is easy to check that $A$ is SPD with condition number $\kappa_{2}(A)\approx 2*10^{6}$. If the $IC(0)$ factorization is computed in exact arithmetic then the entry $(5,5)$ of $L$ is zero (B1 breakdown). The look-ahead strategy reveals this at the third step, thus reducing the work performed before breakdown is detected.
A consequence of look-ahead is that, through the early detection of B1 breakdowns and taking action to prevent such breakdowns, B3 breakdowns are indirectly prevented. In our numerical experiments on problems coming from real applications, all breakdowns when using fp16 arithmetic were of type B1 when look-ahead was incorporated. However, B2 and B3 breakdowns remain possible. Consider the following well-conditioned SPD matrix and its complete Cholesky factor (to three decimal places)
@@
In fp16 arithmetic, the $(5,4)$ entry overflows (B3 breakdown). This happens even with look-ahead because the first three entries in row 5 of $A$ are zero and so the $(5,4)$ entry is not updated until after column 4 has been computed. In this example, the $(4,4)$ entry before its square root is taken is greater than $\tau_{u}$ (it is equal to $7*10^{-5}$) and so there is no B1 breakdown in column 4.
@@ 3.3 Global shifting to handle breakdown
Once potential breakdown has been detected and the factorization halted, a common approach is to modify all the diagonal entries by selecting $\alpha>0$, replacing the scaled matrix $\widehat{A}$ by $\widehat{A}+\alpha I$ and restarting the factorization. There is always an $\alpha^{\ast}$ such that for all $\alpha\geq\alpha^{\ast}$ the IC factorization of $\widehat{A}+\alpha I$ exists [17]. In practice, $\alpha^{\ast}$ is unlikely to be known a priori and it may be necessary to restart the factorization a number of times with ever larger shifts. Algorithm 3 summarizes the global shifting strategy for a SPD matrix held in precision $u$ and for which an IC factorization in precision $u_{l}\geq u$ is wanted. After each unsuccessful factorization attempt, the shift is doubled [15]. Here $fl_{l}(\widehat{A})$ denotes converting the matrix $\widehat{A}$ from precision $u$ to precision $u_{l}$. In our experiments, we use $\alpha_{0}=0$ and $\alpha_{S}=10^{-3}$. The precise choice of the shift is not critical but it should not be unnecessarily large as this may result in the computed factors providing poor quality preconditioners. Note that more sophisticated strategies for changing the shift, which also allow the possibility for a shift to be decreased, are possible [16]. These have been developed for fp64 arithmetic. Our initial experiments suggest it is less clear that there are significant benefits of doing this when using fp16 arithmetic so in our experiments we only report results for using the simple shift doubling strategy.
@@
Input: SPD matrix $A$ in precision $u$, diagonal scaling matrix $S$ and initial shift $\alpha_{S}>0$
@@ 3.4 Local modifications to prevent breakdown
The next strategy is based on seeking to guarantee the factorization exists by bounding the off-diagonal entries in $L$. So-called modified Cholesky factorization schemes have been widely used in nonlinear optimization to compute Newton-like directions. Given a symmetric (and possibly indefinite) $A$, a modified Cholesky algorithm factorizes $A+A_{E}$, where $A_{E}$ is termed the correction matrix. The objectives are to compute the correction at minimal additional cost and to ensure $A+A_{E}$ is SPD and well-conditioned and close to $A$. A stable approach for dense matrices was originally proposed by Gill and Murray [18] and was subsequently refined by Gill, Murray and Wright (GMW) [19] and others [20, 21, 22, 23].
@@
$l_{kk}=\max{\left\{l_{kk},\ \left(\frac{l_{max}}{\beta}\right)^{2}\right\}},$ (3.1)
where $l_{max}$ is given by (2.1). If $(l_{max}/{\beta})^{2}$ overflows (this can be safely checked) then the diagonal entry cannot be modified in this way. We call this a B4 breakdown. If despite the local modification potential breakdown is detected then the factorization is terminated and restarted using a global shift. As the following result shows, the GMW$(\beta)$ strategy limits the size of the off-diagonal entries in $L$ and, for $\beta$ sufficiently small, it prevents B3 breakdown.
@@
Let the matrix $A$ be sparse and SPD. Assume that, using the GMW($\beta$) strategy, columns 1 to $j-1$ columns of the IC factor $L$ have been successfully computed in fp16 arithmetic. For $i\geq j$ let $nz(i)$ denote the number of nonzero entries in $L_{i,1:j-1}$. If $|a_{ij}|+\min(nz(i),nz(j))\beta^{2}\leq x_{max}$ for all $(i,j)\in{\mathcal{S}}\{L\}$ ($x_{max}$ is the largest finite number represented in fp16) then B3 breakdown cannot occur in the $j$-th step.
@@
For $i>j$, entry $(i,j)$ of $L$ is given by
$l_{ij}=\frac{1}{\sqrt{l_{jj}}}\left(a_{ij}-\sum_{\begin{subarray}{c}k=1\\ l_{ik}\neq 0\\ l_{jk}\neq 0\end{subarray}}^{j-1}l_{ik}l_{jk}\right)=\frac{\widetilde{l}_{ij}}{\sqrt{l_{jj}}}.$
From (3.1), the offdiagonal entries in the first $j-1$ columns of $L$ satisfy
$\frac{|l_{ik}|}{\sqrt{l_{kk}}}\leq\frac{|l_{ik}|\beta}{l_{max}}\leq\beta,\quad 1\leq k\leq j-1,\ i>k.$
Consequently,
@@
and B3 breakdown does not occur. ∎
Rules to determine the parameter $\beta$ for the complete Cholesky factorization of dense matrices (in double precision) are discussed in [19, 18] but for sparse incomplete factorizations in fp16 arithmetic such sophisticated rules are not applicable. Provided $\beta$ is not very small, its value is not critical to the quality of the preconditioner. Given $\beta$, Algorithm 4 incorporates the use of the GMW($\beta$) strategy within the incomplete factorization. Note that the local modifications are not combined with look-ahead.
@@
Output: Either $flag=0$ and $A\approx L^{T}L$ or $flag>0$ (breakdown detected)
1:Initialize $l_{ij}=a_{ij}$ for all $(i,j)\in{\mathcal{S}}\{L\}$
@@
3:for $k=1:n$ donormal-$\triangleright$ Start of $k$-th major step
4:     if ${(l_{max}}/{\beta)^{2}}$ does not overflow thennormal-$\triangleright$ $l_{max}$ is largest off-diagonal entry (2.1)
@@ 3.5 Recovering double precision accuracy
Having computed an incomplete factorization in low precision, we seek to recover double precision accuracy in the computed solution (although in many applications much less accuracy may be sufficient and all that is justified but the accuracy in the data). In their work on using mixed precision for solving general linear systems, Carson and Higham [2] introduce a variant of iterative refinement that uses GMRES preconditioned by the LU factors of the matrix to solve the correction equation. Carson and Higham use two precisions and this was later extended to allow up to five precisions [3, 4]; see Algorithm 5. In the SPD case, a natural choice is the conjugate gradient (CG) method in place of GMRES. However, the supporting rounding error analysis relies on the backward stability of GMRES and preconditioned CG is not guaranteed to be backward stable [24]. Note, however, that mixed precision results presented in [5, 6] suggest that in practice CG-IR often works as well as GMRES-IR.
Algorithm 5 GMRES-based iterative refinement using five precisions (IC-Krylov-IR)
@@
5:     Use preconditioned GMRES to solve $Ad_{i}=r_{i}$, with preconditioning and products with $A$ performed in precision $u_{p}$; store $d_{i}$ in precision $u$
@@ 4 Numerical experiments
We follow a number of others working on the development of numerical linear algebra algorithms in mixed precision in performing experiments that aim to explore the feasibility of the ideas by simulating half precision (see, for example, [7, 5, 25, 26]). We want the option to experiment with sparse problems that may be too large for MATLAB and have chosen to develop our software in Fortran. We use the NAG compiler (Version 7.1, Build 7118), which is the only Fortran compiler that currently supports the use of fp16. The NAG documentation states that their half precision implementation conforms to the IEEE standard. In addition, using the -roundhreal option, all half-precision operations are rounded to half precision, both at compile time and runtime. Our numerical experiments are performed on a Windows 11-Pro-based machine with an Intel(R) Core(TM) i5-10505 CPU processor (3.20GHz).
Our test set of SPD matrices is given in Table 1. This set was used in our earlier study [6]. The problems come from a variety of application areas and are of different sizes and densities. As we expect that successfully using fp16 arithmetic will be most challenging for ill-conditioned problems, the problems were chosen because they all have a large estimated condition number (in the range $10^{7}-10^{16}$). Many are initially poorly scaled and some contain entries that overflow in fp16 and thus prescaling of $A$ is essential. The right-hand side vector $b$ is constructed by setting the solution $x$ to be the vector of 1’s. In Table 1, we also report the number of entries in the “scaled and squeezed” matrix $A_{l}$ (see Line 2 of Algorithm 3). The squeezing discards all entries of the scaled matrix with absolute value less than $\tau_{u}=10^{-5}$. We see that this can lead to the loss of a significant number of entries.
Table 1: Statistics for our ill-conditioned test examples. $nnz(A)$ denotes the number of entries in the lower triangular part of $A$. $normA$ and $normb$ are the infinity norms of $A$ and $b$. $cond2$ is a computed estimate of the condition number of $A$. $A_{l}$ is the number of entries in the lower triangular part of the matrix after scaling and squeezing.
@@
Our results are for the level-based incomplete Cholesky factorization preconditioner $IC(\ell)$ with $\ell=2$ and 3 [27]. We refer to the $IC(\ell)$ factorizations computed using half and double precision arithmetic as fp16-$IC(\ell)$ and fp64-$IC(\ell)$, respectively. The key difference between the two versions is that for the former, during the incomplete factorization, we incorporate the safe checks for the scaling and update operations; for the fp64 version, only tests for problem B1 are performed (problems B2 and B3 are not encountered in our double precision experiments). The solves with $L$ and $L^{T}$ require the $L$ factor to be in double precision. This can be done by casting the data into double precision and making an explicit copy of $L$; this negates the important benefit that half precision offers of reducing memory requirements. Alternatively, the entries can be cast on the fly. This is straightforward to incorporate into a serial triangular solve routine, and only requires a temporary double precision array of length $n$.
@@
We use two precisions: $u_{\ell}=u_{16}$ for the incomplete factorization and $u_{r}=u_{g}=u_{p}=u_{64}$, where $u_{16}$ and $u_{64}$ denote the unit roundoffs in fp16 and fp64 arithmetic, respectively. That is, we aim to achieve double precision accuracy in the computed solution. The iterative refinement terminates when the normwise backward error for the computed solution satisfies
$res=\frac{\|b-Ax\|_{\infty}}{\|A\|_{\infty}\|x\|_{\infty}+\|b\|_{\infty}}\leq\delta=10^{3}\times u_{64}.$ (4.1)
@@
The implementation of GMRES is taken from the HSL software library [28], and its convergence tolerance is set to $u_{64}^{1/4}$ (see [6] for an explanation of this choice); for each application of GMRES the limit on the number of iterations is 1000.
@@
Our first experiment looks at the effects of incorporating look-ahead. Table 2 presents results with no look-ahead and with look-ahead. Here we only include the test problems for which look-ahead had an effect. We can make a number of observations. Incorporating look-ahead can improve robustness, particularly when using double precision. Without look-ahead, in fp64 arithmetic there can be very large growth in the size of some entries in the factors and this goes undetected (no B1 problems are found with $\tau_{u}=10^{-20}$). With look-ahead, growth is potentially possible but it did not occur in our tests on ill conditioned problems. In fp16, look-ahead can replace B3 breakdown by B1 breakdown (e.g., HB/bcsstk24). Even if there are no B3 breakdowns, look-ahead can lead to a larger number of B1 breakdowns being flagged (Boeing/msc01050) and hence a larger number of restarts, a larger shift and, consequently, a higher GMRES iteration count. Note that the iteration counts for $IC(3)$ are not guaranteed to be smaller than for $IC(2)$ (although they generally are).
@@
Table 2: Results for IC-GMRES-IR using fp16-$IC(\ell)$ and fp64-$IC(\ell)$ preconditioners ($\ell=2$, 3) with no look-ahead and with look-ahead. $its$ is the total number of GMRES iterations and $(n1,n2)$ are the numbers of times B1 and B3 problems are detected ($n2$ is nonzero only for no look-ahead). $>1000$ indicates the requested accuracy was not obtained within the GMRES iteration limit. ‡ indicates failure to compute useful factors because of enormous growth in the entries.
@@
Cylshell/s3rmt3m1 102 (0, 0) 102 (0) $>1000$ (0, 0) 426 (1)
Boeing/ct20stif $>1000$ (0, 0) 1940 (2) 1332 (0, 0) 1368 (1)

### v4

@@ 3.2 Preventing breakdown by incorporating look-ahead
Initially, $l_{jj}=a_{jj}$ and at each stage of the factorization a positive (or zero) term is subtracted from it so that $l_{jj}$ either decreases or remains the same on each major step $k$. Thus, to detect potential B1 breakdown as early as possible, look-ahead can be used whereby, at each step $k$, the remaining diagonal entries $l_{jj}$ ($j>k$) are updated (using safe operations) and tested. For the right-looking Algorithm 1, it is straightforward to incorporate testing but, for some IC variants, it may be necessary to hold a copy of the diagonal entries of the factor. Look-ahead is employed in some well-known fp64 arithmetic implementations of IC factorizations e.g., [20, 21]. Algorithm 2 is a modified version of Algorithm 1 that includes checks for breakdown. The safe test for B3 breakdown can be modified so that, ahead of the loop at Line 13, a check is made that the entry of maximum magnitude in column $k$ (that is, $l_{kmax}$ from (2.1)) is less than $(x_{max})^{1/2}$. If it is not, $flag=3$ is returned. Otherwise, the multiplication of $l_{ik}$ and $l_{jk}$ is safe and, at Line 17, only the subtraction needs to be checked. If in place of the safe tests, IEEE exception handling is used then the IEEE overflow flag should be tested at the end of each major loop (that is, between Lines 20 and 21).
Algorithm 2 Right-looking IC factorization with safe checks for breakdown
@@
Output: Either $flag=0$ and $A\approx LL^{T}$ or $flag>0$ (breakdown detected)
1:$l_{ij}=a_{ij}$ for all $(i,j)\in{\mathcal{S}}\{L\}$
2:$flag=0$
@@
4:for $k=1:n$ $\triangleright$ Start of $k$-th major step
@@
6:     if  $l_{kk}\geq 1$ or $l_{kk}\geq l_{max}/x_{max}$ $\triangleright$ $l_{max}$ is largest off-diagonal entry (2.1)
7:         for  $i\in\{i>k\,|\,(i,k)\in{\mathcal{S}}\{L\}\}$ do
@@
9:         end $\triangleright$ Column $k$ of $L$ has been computed
@@
11:         $flag=2$ and $\triangleright$ B2 breakdown
@@
13:     for  $j\in\{j>k\,|\,(j,k)\in{\mathcal{S}}\{L\}\}$ do $\triangleright$ Update columns $j>k$ of $L$
14:         for $i\in\{i\geq j\,|\,(i,j)\in{\mathcal{S}}\{L\}\}$ do
15:              Test entry $(i,j)$ can be updated safely $\triangleright$ Use Algorithm 2.2 of [6]
@@
18:         end for
19:         if $l_{ii}<\tau_{u}$ then $flag=1$ and return $\triangleright$ B1 breakdown
20:     end for
@@
To see the usefulness of look-ahead, consider the following matrix
@@
It is easy to check that $A$ is SPD with condition number $\kappa_{2}(A)\approx 2\times 10^{6}$. If the $IC(0)$ factorization is computed in exact arithmetic then the entry $(5,5)$ of $L$ is zero (B1 breakdown). The look-ahead strategy reveals this at the third step, thus reducing the work performed before breakdown is detected.
A consequence of look-ahead is that, through the early detection of B1 breakdowns and taking action to prevent such breakdowns, B3 breakdowns are indirectly prevented. In our numerical experiments on problems coming from real applications, all breakdowns when using fp16 arithmetic were of type B1 when look-ahead was incorporated. However, B2 and B3 breakdowns remain possible. Consider the following well-conditioned SPD matrix and its complete Cholesky factor (three decimal places):
@@
In fp16 arithmetic, the $(5,4)$ entry overflows (B3 breakdown). This happens even with look-ahead because the first three entries in row 5 of $A$ are zero and so the $(5,4)$ entry is not updated until after column 4 has been computed. In this example, the $(4,4)$ entry before its square root is taken is greater than $\tau_{u}$ (it is equal to $7\times 10^{-5}$) and so there is no B1 breakdown in column 4.
@@ 3.3 Global shifting to handle breakdown
Once potential breakdown has been detected (either through the B1-B3 tests or using IEEE exception handling) and the factorization halted, a common approach is to modify all the diagonal entries by selecting $\alpha>0$, replacing the scaled matrix $\widehat{A}$ by $\widehat{A}+\alpha I$ and restarting the factorization. In exact arithmetic, there is always an $\alpha^{\ast}$ such that for all $\alpha\geq\alpha^{\ast}$ the IC factorization of $\widehat{A}+\alpha I$ exists [22]. In practice, $\alpha^{\ast}$ is unlikely to be known a priori and it may be necessary to restart the factorization a number of times with ever larger shifts. If $\alpha$ was to reach $x_{max}$ in the working precision $u$ then there would have been large growth in the factor entries and it would be necessary to restart using higher precision. However, this would be unlikely to result in a useful preconditioner and $\alpha$ growing in this way was not observed in any of our tests (including highly ill-conditioned examples). Algorithm 3 summarizes the global shifting strategy for a SPD matrix held in precision $u$ and for which an IC factorization in precision $u_{l}\geq u$ is wanted. After each unsuccessful factorization attempt, the shift is doubled [20]. Here $A_{l}=fl_{l}(\widehat{A})$ denotes converting the matrix $\widehat{A}$ from precision $u$ to precision $u_{l}$. In practice, it is unnecessary to explicitly hold $A_{l}$. Instead, entries of $\widehat{A}$ are cast to precision $u_{l}$ on the fly as needed.
Based on our experience with a range of problems, in our reported tests, the initial shift is taken to be $\alpha_{S}=10^{-3}$. The precise choice of the shift is not critical but it should not be unnecessarily large as this may result in the computed factors providing poor quality preconditioners. Note that more sophisticated strategies for changing the shift, which also allow the possibility for a shift to be decreased, are possible [21]. These have been developed for fp64 arithmetic. Our initial experiments suggest it is less clear that there are significant benefits of doing this when using fp16 arithmetic so in our experiments we only report results for using the simple shift doubling strategy.
@@
Input: SPD matrix $A$ in precision $u$, diagonal scaling matrix $S$, a target sparsity pattern ${\mathcal{S}}\{L\}$, and initial shift $\alpha_{S}>0$
@@ 3.4 Local modifications to prevent breakdown
The next strategy is based on seeking to guarantee that the factorization exists by bounding the off-diagonal entries in $L$. So-called modified Cholesky factorization schemes have been widely used in nonlinear optimization to compute Newton-like directions. Given a symmetric (and possibly indefinite) $A$, a modified Cholesky algorithm factorizes $A+A_{E}$, where $A_{E}$ is termed the correction matrix. The objectives are to compute the correction at minimal additional cost and to ensure $A+A_{E}$ is SPD and well-conditioned and close to $A$. A stable approach for dense matrices was originally proposed by Gill and Murray [23] and was subsequently refined and used by Gill, Murray and Wright (GMW) [24] and others [25, 26, 27, 28, 29].
@@
$l_{kk}=\max{\left\{l_{kk},\ \left(\frac{{l_{kmax}}}{\beta}\right)^{2}\right\}},$ (3.1)
where $l_{kmax}$ is given by (2.1). If $(l_{kmax}/{\beta})^{2}$ overflows (this can be safely checked) then the diagonal entry cannot be modified in this way. We call this a B4 breakdown. If despite the local modification potential breakdown is detected then the factorization is terminated and restarted using a global shift. As the following result shows, the GMW$(\beta)$ strategy limits the size of the off-diagonal entries in $L$ and, for $\beta$ sufficiently small, it prevents B3 breakdown.
@@
Let the matrix $A$ be sparse and SPD. Assume that, using the GMW($\beta$) strategy, columns 1 to $j-1$ columns of the IC factor $L$ have been successfully computed in fp16 arithmetic. For $i\geq j$ let $nz(i)$ denote the number of nonzero entries in $L_{i,1:j-1}$. If
$|a_{ij}|+\min(nz(i),nz(j))\beta^{2}\leq x_{max}\quad\mbox{for all}\quad(i,j)\in{\mathcal{S}}\{L\},$ (3.2)
where $x_{max}$ is the largest finite number represented in fp16, then B3 breakdown cannot occur in the $j$-th step.
@@
From (3.1), the off-diagonal entries in the first $j-1$ columns of $L$ satisfy
$\frac{|l_{ik}|}{(l_{kk})^{1/2}}\leq\frac{|l_{ik}|\beta}{l_{kmax}}\leq\beta,\quad 1\leq k\leq j-1,\ i>k.$
For any $i>j$ we have
$l_{ij}=\frac{1}{(l_{jj})^{1/2}}\left(a_{ij}-\sum_{k=1}^{j-1}l_{ik}l_{jk}\right):=\frac{\widetilde{l}_{ij}}{(l_{jj})^{1/2}}.$ (3.3)
To avoid breakdown, we require $|\widetilde{l}_{ij}|\leq x_{max}$. From (3.2) and (3.3),
@@
and hence B3 breakdown does not occur.
Rules to determine the parameter $\beta$ for the complete Cholesky factorization of dense matrices (in double precision) are discussed in [24, 23] but for sparse incomplete factorizations in fp16 arithmetic such sophisticated rules are not applicable. Provided $\beta$ is not very small, its value is not critical to the quality of the preconditioner. Given $\beta$, Algorithm 4 incorporates the use of the GMW($\beta$) strategy within the incomplete factorization. Note that the local modifications are not combined with look-ahead. As in Algorithm 2, the cost of checking for breakdowns is small. The most expensive step is computing $l_{max}$.
@@
Output: Either $flag=0$ and $A\approx LL^{T}$ or $flag>0$ (breakdown detected)
1:$l_{ij}=a_{ij}$ for all $(i,j)\in{\mathcal{S}}\{L\}$
@@
3:for $k=1:n$ $\triangleright$ Start of $k$-th major step
4:     if ${(l_{max}}/{\beta)^{2}}$ does not overflow $\triangleright$ $l_{max}$ is largest off-diagonal entry (2.1)
@@ 3.5 Recovering double precision accuracy
Having computed an incomplete factorization in low precision, we seek to recover (close to) double precision accuracy in the final solution (although in many applications much less accuracy may be sufficient and may be all that is justified by the accuracy in the data). In their work on using mixed precision for solving general linear systems, Carson and Higham [2] introduce a variant of iterative refinement that uses GMRES preconditioned by the low precision LU factors of the matrix to solve the correction equation (GMRES-IR). Carson and Higham employ two precisions. This was later extended to three precisions and then to five precisions [3, 4]; see Algorithm 5, where we use a generic Krylov solver and the low precision incomplete factors. Here $u$ is the working precision. In the three-precision variant [4], $u_{p}=u$, $u_{p}=u^{2}$ and $u_{l}\geq u\geq u_{r}$ and typical combinations include $(u_{l},u,u_{r})=(u_{16},u_{32},u_{64})$ or $(u_{16},u_{64},u_{64})$. Section 3.4 of [3] discusses meaningful combinations of the five precisions. They must satisfy $u^{2}\leq u_{r}\leq u\leq u_{l}$, $u_{p}\leq u_{g}$ and $u_{p}<u_{l}$. The use of more than two precisions has the potential to solve problems that are less well conditioned in less time and using less memory.
In the SPD case, a natural choice is to select the conjugate gradient (CG) method to be the Krylov solver. The supporting rounding error analysis for GMRES-IR relies on the backward stability of GMRES and preconditioned CG is not guaranteed to be backward stable [30]. Nevertheless, mixed precision results presented in [5] (using MATLAB code and relatively small test examples) suggest that in practice CG-IR can perform as well as GMRES-IR. In our earlier paper [6], we experiment with IC-CG-IR and compare it with IC-GMRES-IR using fp16 and fp64 arithemtic. While there is little to choose between them when run on well-conditioned SPD problems, for highly ill-conditioned examples, IC-CG-IR often (but not always) requires a greater number of iterations to obtain double precision accuracy.
Algorithm 5 IC-Krylov-IR: Krylov solver iterative refinement using five precisions
@@
5:     Use a preconditioned Krylov solver to solve $Ad_{i}=r_{i}$ at precision $u_{g}$, with preconditioning and products with $A$ in precision $u_{p}$; store $d_{i}$ in precision $u$
@@ 4 Numerical experiments
We follow a number of others working on the development of numerical linear algebra algorithms in mixed precision in performing experiments that aim to explore the feasibility of the ideas by using half precision (see, for example, [8, 5, 31, 32]). We want the option to experiment with sparse problems that may be too large for MATLAB and have chosen to develop our software in Fortran. We use the NAG compiler (Version 7.1, Build 7118). As far as we know, it is the only multi-platform Fortran compiler that currently fully supports the use of fp16 arithmetic and conforms to the IEEE standard. In addition, using the -roundhreal option, all half-precision operations are rounded to half precision, both at compile time and runtime. Our numerical experiments are performed on a Windows 11-Pro-based machine with an Intel(R) Core(TM) i5-10505 CPU processor (3.20 GHz).
Our test set of SPD matrices is given in Table 1. This set was used in our earlier study [6]. For consistency with that study, we do not reorder the matrix $A$. The problems come from a variety of application areas and are of different sizes and densities. As we expect that successfully using fp16 arithmetic will be most challenging for ill-conditioned problems, the problems were chosen because they all have a large estimated condition number (in the range $10^{7}-10^{16}$). Many are initially poorly scaled and some contain entries that overflow in fp16 and thus prescaling of $A$ is essential. The right-hand side vector $b$ is constructed by setting the solution $x$ to be the vector of 1’s. In Table 1, we also report the number of entries in the “scaled and squeezed” matrix $A_{l}$ (see Line 2 of Algorithm 3). The squeezing discards all entries of the scaled matrix with absolute value less than $\tau_{u}=10^{-5}$. We see that this can lead to the loss of a significant number of entries.
Table 1: Statistics for our ill-conditioned test examples. $nnz(A)$ denotes the number of entries in the lower triangular part of $A$. $normA$ and $normb$ are the infinity norms of $A$ and $b$. $cond2$ is a computed estimate of the condition number of $A$ in the 2-norm. $nnz(A_{l})$ is the number of entries in the lower triangular part of the matrix after scaling and squeezing.
@@
Our results are for the level-based incomplete Cholesky factorization preconditioner $IC(\ell)$ with $\ell=2$ and 3 [33]. The number of entries in the incomplete factor $L$ increases as the parameter $\ell$ increases. $IC(0)$ is a very simple preconditioner in which $L$ has the same sparsity pattern as $A$. In practice, using very small $\ell$ may be sufficient for solving well-conditioned problems but, as shown in [6], the resulting preconditioner is often not powerful enough to successfully tackle ill-conditioned examples (particularly when computed using fp16 arithmetic). We refer to the $IC(\ell)$ factorizations computed using half and double precision arithmetic as fp16-$IC(\ell)$ and fp64-$IC(\ell)$, respectively. The key difference between the two versions is that for the former, during the incomplete factorization, we incorporate the safe checks for the scaling and update operations; for the fp64 version, tests for B1 breakdown are performed (B2 and B3 breakdowns were not encountered in our double precision experiments). The solves with $L$ and $L^{T}$ employ the $L$ factor in double precision. This can be done by casting the data into double precision and making an explicit copy of $L$; this negates the important benefit that half precision offers of reducing memory requirements. Alternatively, the entries can be cast on the fly. This is straightforward to incorporate into a serial triangular solve routine, and only requires a temporary double precision array of length $n$. This is done in our software.
@@
We use two precisions: $u_{\ell}=u_{16}$ for the incomplete factorization and $u_{r}=u_{g}=u_{p}=u=u_{64}$, where $u_{16}$ and $u_{64}$ denote the unit roundoffs in fp16 and fp64 arithmetic, respectively. That is, we aim to achieve double precision accuracy in the computed solution. The iterative refinement terminates when the normwise backward error for the computed solution satisfies
$res=\frac{\|b-Ax\|_{\infty}}{\|A\|_{\infty}\|x\|_{\infty}+\|b\|_{\infty}}\leq\delta=10^{3}\times u_{64}.$
@@
The implementation of GMRES is taken from the HSL software library [34] (MI24 is a Fortran MGS-GMRES implementation), and its convergence tolerance is set to $u_{64}^{1/4}$ (see [6] for an explanation of this choice); for each application of GMRES (Step 5 of Algorithm 5) the limit on the number of iterations is 1000. Restarting is not used. In the tables of results, NC denotes that this limit has been exceeded without the GMRES convergence tolerance being achieved.
@@
Our first experiment looks at the effects of incorporating look-ahead. Table 2 presents results with no look-ahead and with look-ahead. Here we only include the test problems for which look-ahead has an effect. We make a number of observations. Incorporating look-ahead can improve robustness, particularly when using fp64 arithmetic. Without look-ahead, in fp64 arithmetic there can be very large growth in the size of some entries in the factors and this goes undetected (no B1 breakdowns occur with $\tau_{u}=10^{-20}$). With look-ahead, growth did not happen in our tests on ill-conditioned problems. In fp16 arithmetic, look-ahead can replace B3 breakdown by B1 breakdown (e.g., HB/bcsstk24). Even if there are no B3 breakdowns, look-ahead can lead to a larger number of B1 breakdowns being flagged (see problem Boeing/msc01050) and hence a larger number of restarts, a larger shift and, consequently, a higher GMRES iteration count. Note that the iteration counts for $IC(3)$ are not guaranteed to be smaller than for $IC(2)$ (although they generally are).
@@
Table 2: Results for IC-GMRES-IR (Algorithm 5) using fp16-$IC(\ell)$ and fp64-$IC(\ell)$ preconditioners ($\ell=2$, 3) with no look-ahead and with look-ahead (Section 3.2). $its$ is the total number of GMRES iterations and $(n1,n2)$ are the numbers of times B1 and B3 breakdowns are detected ($n2$ is nonzero only for fp16-$IC(\ell)$ with no look-ahead). NC indicates that on an inner iteration the requested GMRES accuracy of $u_{64}^{1/4}$ was not achieved within the limit of 1000 iterations. ‡ indicates failure to compute useful factors because of enormous growth in the entries.
@@
Cylshell/s3rmt3m1 102 (0, 0) 102 (0) NC (0, 0) 426 (1)
Boeing/ct20stif NC (0, 0) 1940 (2) 1332 (0, 0) 1368 (1)